OpenStudy (anonymous):

Write the complex number in the form a + bi. 5/2(cos 150° + i sin 150°)

OpenStudy (jdoe0001):

\(\bf r[cos(\theta)+i\ sin(\theta)]\to (r,\theta)\) what do you think the coordinates will be there? that is, what's "r" and what is the angle?

OpenStudy (anonymous):

ooh so 5/2+150i ?

OpenStudy (jdoe0001):

anyhow... once you find that, keep in mind that \(\large \begin{cases} cos(\theta)=\frac{x}{r}\to rcos(\theta)={\color{blue}{ x}} \\ \quad \\ sin(\theta)=\frac{y}{r}\to rsin(\theta)={\color{blue}{ y}} \end{cases}\)

OpenStudy (jdoe0001):

and that'd give the (x,y) rectangular coordinates...which can be written as x + yi

OpenStudy (jdoe0001):

and yes... the r = 5/2 and the angle is 150 degrees so get "x" and "y" :)

OpenStudy (ybarrap):

 $$ \cfrac{5}{2}\left (\cos 150 + i\sin 150\right )=\cfrac{5}{2}\left (\cfrac{-\sqrt 3}{2}+i\cfrac{1}{2}\right )\\ =-\cfrac{5\sqrt 3}{4}+\cfrac{5}{4}=5\cfrac{1-\sqrt 3}{4} $$ Does this make sense?